# UMGAD: Unsupervised Multiplex Graph Anomaly Detection

### v2

@@ Abstract
Graph anomaly detection (GAD) is a critical task in graph machine learning, with the primary objective of identifying anomalous nodes that deviate significantly from the majority. This task is widely applied in various real-world scenarios, including fraud detection and social network analysis. However, existing GAD methods still face two major challenges: (1) They are often limited to detecting anomalies in single-type interaction graphs and struggle with multiple interaction types in multiplex heterogeneous graphs. (2) In unsupervised scenarios, selecting appropriate anomaly score thresholds remains a significant challenge for accurate anomaly detection. To address the above challenges, we propose a novel Unsupervised Multiplex Graph Anomaly Detection method, named UMGAD. We first learn multi-relational correlations among nodes in multiplex heterogeneous graphs and capture anomaly information during node attribute and structure reconstruction through graph-masked autoencoder (GMAE). Then, to further extract abnormal information, we generate attribute-level and subgraph-level augmented-view graphs respectively, and perform attribute and structure reconstruction through GMAE. Finally, We learn to optimize node attributes and structural features through contrastive learning between original-view and augmented-view graphs to improve the model’s ability to capture anomalies. Meanwhile, we also propose a new anomaly score threshold selection strategy, which allows the model to be independent of ground truth information in real unsupervised scenarios. Extensive experiments on four datasets show that our UMGAD significantly outperforms state-of-the-art methods, achieving average improvements of 13.48% in AUC and 11.68% in Macro-F1 across all datasets. The source code of our model is available at https://anonymous.4open.science/r/UMGAD-1F3B.
@@ I Introduction
Challenges. Existing UGAD approaches, such as CL-based and GAE-based methods, have achieved promising results. Nevertheless, the UGAD task still faces the following two major challenges: (1) Most existing methods focus solely on non-multiplex heterogeneous graphs, while real-world graphs are often multiplex that typically include multiple types of interactions. For example, there exists viewing, carting, and buying relations between users and items in e-commerce networks, or users’ different comments and rating scores on items in review networks. These interactions result in complex structures, known as multiplex heterogeneous graphs. Anomaly detection in multiplex heterogeneous graphs is challenging because the interacting multiple relations are extremely complex and different types of relations have different effects on anomaly detection. (2) In real unsupervised scenarios, most of existing aspects face difficulties in selecting anomaly score thresholds because the number of anomalies is unknown. Existing models typically employ two approaches to select the anomaly score threshold: 1) by selecting the threshold based on the known number of anomalies when ground truth information in the test set is available, and 2) by choosing the threshold that yields the best model performance. However, these methods are not suitable when the number of anomalies and their labels are unknown, making it unreasonable to rely on these approaches for threshold selection in the real unsupervised scenario with no ground truth information from the test set.
@@
Presented Work. Recognizing the above challenges, we focus on exploring the important impact of different interactive relations on node representation learning and extracting anomaly information through graph reconstructions for multiplex heterogeneous graphs. To this end, this paper proposes a novel unsupervised multiplex graph anomaly detection method, named UMGAD. We first learn multi-relational correlations among nodes in multiplex heterogeneous graphs and capture anomaly information during node attribute and structure reconstruction in the original-view graph. Then, to further extract abnormal information, we generate attribute-level and subgraph-level augmented-view graphs respectively, and perform attribute and structure reconstruction in the augmented-view graph. Innovatively, we coordinate the fusion of GMAEs with multiple masking mechanisms across multiplex graphs to reconstruct node attributes, network structures, and subgraphs in the original and augmented graphs. Furthermore, we learn to optimize node attributes and structural features through contrastive learning between original-view and augmented-view graphs to improve the model’s ability to capture anomalies. Meanwhile, we also propose a new anomaly score threshold selection strategy, which allows the model to be independent of the ground truth in real unsupervised scenarios. Extensive experiments on datasets with synthetic and real anomalies have demonstrated the effectiveness and efficiency of UMGAD compared with state-of-the-art UGAD methods.
@@
  • We propose a novel unsupervised multiplex graph anomaly detection method named UMGAD that emphasizes the importance of constructing and exploiting different interactions between nodes and addresses the problem of unsupervised anomaly detection in multiplex heterogeneous graphs.
@@
  • We conduct extensive experiments on datasets with both synthetic and real anomalies to demonstrate the superiority and efficiency of our proposed method. Experiment results show that UMGAD achieves 13.48%, and 11.68% average improvement in terms of AUC and macro-F1 across four datasets compared to SOTA baselines.
@@ II Related Work
UGAD has attracted increasing attention in recent years, many researchers focus on how to use self-supervised signals to improve the model’s ability to detect anomalies in unsupervised scenarios. Early traditional UGAD methods, such as Radar [27] and ResGCN [34], which capture anomalous features by describing the residuals of attribute information and their consistency with network information, are concise and effective. With the rapid development of GNNs, it has also been naturally introduced into the field of anomaly detection, and recent researchers have been trying to develop UGAD models based on GNNs. We broadly categorize GNN-based models into four types: MPI methods, CL-based methods, GAE-based methods and MV methods.
@@
MPI methods. It is well known that most of the GNN-based methods exploit the message passing mechanism [35] of GNN to learn node attributes by aggregating the neighborhood information of nodes. However, due to the feature inconsistency between normal and anomalous nodes, the original message passing mechanism may impair the model’s ability to extract the anomalous information. Luo et al. [28] enhance the message passing mechanism through community segmentation and structure learning, obtaining node attributes by introducing the community structure. Ber et al. [36] propose RAND, a reinforcement learning-based method for unsupervised graph anomaly detection, which enriches neighbor pools, evaluates reliability, selects the best subset, and amplifies messages from reliable neighbors. Qiao et al. [29] propose TAM, a truncated affinity maximization method that maximizes local affinity with neighbors and performs node characterization on a truncated graph instead of the original graph.
@@
CL-based methods. As a common method for mining self-supervised signals [37, 38, 39], contrastive learning is also introduced to the UGAD task [40] to capture feature inconsistencies between normal and anomalous nodes. CoLA [41] introduces a contrastive instance pair sampling method that leverages local network information to embed high-dimensional attributes and local structures. Zhang et al. [42] propose Sub-CR, a self-supervised model combining multi-view contrastive learning and attribute reconstruction for anomaly detection on attribute networks. GRADATE [31] extends contrastive learning to the subgraph level with node-subgraph and subgraph-subgraph contrastive learning, effectively identifying complex structural anomalies. Huang et al. [32] integrate variance-based and attribute reconstruction models for balanced outlier detection.
@@
Figure 1: The overview of the proposed UMGAD. There are three interaction types in the multiplex heterogeneous graph: U-P-U links users reviewing at least one product, U-S-U link users having at least one identical star rating within a week, U-V-U links linguistically similar users.
@@
GAE-based methods. GAE-based methods [43] are one of the mainstay methods to detect anomalies by reconstructing the node attributes and structural features in the graphs. Ding et al. [44] leverage GCN to model graph topology and node attributes for learning node embeddings. Roy et al. [45] propose GAD-NR, which reconstructs a node’s entire neighborhood based on its representation, encompassing local structure, self-properties, and neighborhood properties. ADA-GAD [33] introduces a two-stage anomaly denoising autoencoder framework. The first stage trains the GAE on denoised graphs, while the second retrains the encoder on original graphs with anomaly distribution regularization to prevent overfitting.
@@
MV methods. Additionally, some recent methods have focused on the GAD task on multiplex graphs, which we categorize as multi-view GAD methods. AnomMAN [46] detects anomalies on multi-view attribute networks by considering attributes and interactions across views, using an attention mechanism to fuse their importance. DualGAD [13] adopts a dual-bootstrapped self-supervised approach to tackle feature-structure inconsistency and imbalanced anomaly signals. It combines subgraph reconstruction with a cluster-guided contrastive learning mechanism, achieving state-of-the-art performance in node-level anomaly detection.
@@ IV-A1 Attribute Reconstruction
where the original node attributes in $\mathcal{X}_{ma}$ are replaced by the [MASK] tokens which are learnable vectors. Masked nodes are selected by utilizing uniform random sampling without replacement, which helps prevent potential bias. We repeat the attribute masking strategy on each relational subgraph for $K$ times to finally generate a set of masked subgraphs $\mathcal{G}_{ma}^{r}=\{\mathcal{G}_{ma}^{r,1},\mathcal{G}_{ma}^{r,2},\ldots,\mathcal{G}_{ma}^{r,K}\}$ for $\mathcal{G}^{r}$. These masked subgraphs are then fed into the GNN-based encoder and decoder and the output of each encoder is:
@@
where $\mathcal{X}_{ma}^{r,k}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the output of the attribute decoder in the $k$-th masking repeat of the $r$-th relational subgraph. $\mathcal{V}_{re}=\mathcal{V}\setminus\mathcal{V}_{ma}$ denotes the set of remaining nodes that are unaffected by masking, $\mathbf{W}_{enc1}^{r,k}\in\mathbb{R}^{f\times d_{h}}$ and $\mathbf{W}_{dec1}^{r,k}\in\mathbb{R}^{d_{h}\times f}$ are the trainable weight matrices. Inside each encoder and decoder, before the vertical lines |||| are the input variables, and after the vertical lines are the important learnable weights. In subsequent formulas, each vertical line means the same.
After processing each decoupled relational subgraph, UMGAD learns the importance (weights) of different relations collaboratively, instead of dealing with each relational subgraph independently. It fuses the effects of these relations on anomaly detection through a set of learnable weight parameters $a^{r}$ learnable weights as follows:
@@ IV-A2 Structure Reconstruction
where $g(v,u)={x_{ms}^{r,k}(v)}^{\mathsf{T}}x_{ms}^{r,k}(u)$ calculates the estimated probability of link between node $v$ and $u$, and then we can obtain the reconstructed adjacent matrix $\tilde{\mathbf{A}}_{O}^{r}$. We introduce negative sampling to train a more generalized model by accumulating all edge prediction probabilities of the unmasked subgraphs in the denominator, making the model robust to external noise. Notably, the standard normalization in Eq. 4 is used to calculate the similarity between the reconstructed attribute features and the original attribute features. The softmax normalization in Eq. 7 is used to evaluate the probability of a link being reconstructed between two nodes.
Similar to Eq. 3, collaboratively considering the importance of different relational subgraphs for structure reconstruction, we aggregate all subgraph structure reconstruction losses together using a set of learnable weight parameters $b^{r}$ as follows:
@@ IV-B2 Subgraph-level Augmented Graph Reconstruction
Subgraph-level augmentation can be considered as a specific combination of node attribute level and structure level augmentation. The reconstructed node attribute matrix $\tilde{\mathcal{X}}_{sa}^{k}$ and structural adjacent matrix $\tilde{\mathbf{A}}_{ss}^{r}$ are shown below, respectively:
@@
Then the subgraph-level augmented attribute and structure reconstruction loss values $\mathcal{L}_{sa}$ and $\mathcal{L}_{ss}$ is defined as follows:
@@
It is worthwhile to emphasize that normal nodes tend to have more similar features than their neighbors, while anomalous nodes exhibit greater feature differences. Through masking and GNN-based reconstruction, normal nodes are more likely to be accurately reconstructed, whereas anomalous nodes are less likely to be. Therefore, we improve the model’s ability to identify anomalies by employing different masking mechanisms with graph autoencoder.
@@ IV-C Dual-view Contrastive Learning
Dual-view contrastive learning is defined between the original view and the augmented view, to learn more representative and intrinsic node embeddings, which helps capture further anomaly information. Specifically, we optimize the node attribute under joint loss by comparing the original-view graph with the attribute-level augmented-view graph and the subgraph-level augmented-view graph, respectively. The node attribute of node $i$ in the original view forms a positive pair with the augmented view and a negative pair with the attribute of node $j$ in the original and augmented views. We use the following loss function to optimize contrastive learning:
@@ IV-D Optimization Objective
Putting it all together, we have the overall loss function in the training stage is as follows:
@@
where {*} denotes the specific original $(O)$, attribute-level augmented $(A\_Aug)$, or subgraph-level augmented $(S\_Aug)$ graph, $||\cdot|{|_{1}}$ denotes the Euclidean norm of a vector, $||\cdot|{|_{2}}$ denotes the L1-norm of the vector, $\tilde{\zeta}_{*}^{r}(i)$ and $\zeta^{r}(i)$ represent the $i$-th row of the specific reconstructed structure matrix $\tilde{\mathbf{A}}_{*}^{r}$ and original structure matrix $\mathbf{A}^{r}$, respectively. Similarly, $\tilde{x}_{*}(i)$ and $x(i)$ are reconstructed attribute vector and original attribute vector of the node $i$, respectively. Therefore, the anomaly score $S(i)$ of node $i$ is the mean value of all $S(i)_{*}$.

### v4

@@ Abstract
Graph anomaly detection (GAD) is a critical task in graph machine learning, with the primary objective of identifying anomalous nodes that deviate significantly from the majority. This task is widely applied in various real-world scenarios, including fraud detection and social network analysis. However, existing GAD methods still face two major challenges: (1) They are often limited to detecting anomalies in single-type interaction graphs and struggle with multiple interaction types in multiplex heterogeneous graphs. (2) In unsupervised scenarios, selecting appropriate anomaly score thresholds remains a significant challenge for accurate anomaly detection. To address the above challenges, we propose a novel Unsupervised Multiplex Graph Anomaly Detection method, named UMGAD. We first learn multi-relational correlations among nodes in multiplex heterogeneous graphs and capture anomaly information during node attribute and structure reconstruction through graph-masked autoencoder (GMAE). Then, to further extract abnormal information, we generate attribute-level and subgraph-level augmented-view graphs, respectively, and perform attribute and structure reconstruction through GMAE. Finally, we learn to optimize node attributes and structural features through contrastive learning between original-view and augmented-view graphs to improve the model’s ability to capture anomalies. Meanwhile, we propose a new anomaly score threshold selection strategy, which allows the model to be independent of ground truth information in real unsupervised scenarios. Extensive experiments on six datasets show that our UMGAD significantly outperforms state-of-the-art methods, achieving average improvements of 12.25% in AUC and 11.29% in Macro-F1 across all datasets.
@@ I Introduction
Challenges. Existing UGAD approaches, such as CL-based and GAE-based methods, have achieved promising results. Nevertheless, the UGAD task still faces the following two major challenges: (1) Most existing methods focus solely on non-multiplex heterogeneous graphs, while real-world graphs are often multiplex that typically include multiple types of interactions. For example, there exist viewing, carting, and buying relations between users and items in e-commerce networks, or users’ different comments and rating scores on items in review networks. These interactions result in complex structures, known as multiplex heterogeneous graphs [34, 35, 36]. Anomaly detection in multiplex heterogeneous graphs is challenging because the interacting multiple relations are extremely complex and different types of relations have different effects on anomaly detection. (2) In real unsupervised scenarios, most of existing aspects face difficulties in selecting anomaly score thresholds because the number of anomalies is unknown. Existing models typically employ two approaches to select the anomaly score threshold: 1) by selecting the threshold based on the known number of anomalies when ground truth information in the test set is available, and 2) by choosing the threshold that yields the best model performance. However, these methods are not suitable when the number of anomalies and their labels are unknown, making it unreasonable to rely on these approaches for threshold selection in the real unsupervised scenario with no ground truth information from the test set.
@@
Presented Work. Recognizing the above challenges, we focus on exploring the important impact of different interactive relations on node representation learning and extracting anomaly information through graph reconstructions for multiplex heterogeneous graphs. To this end, this paper proposes a novel unsupervised multiplex graph anomaly detection method, named UMGAD. We first learn multi-relational correlations among nodes in multiplex heterogeneous graphs and capture anomaly information during node attribute and structure reconstruction in the original-view graph. Then, to further extract abnormal information, we generate attribute-level and subgraph-level augmented-view graphs respectively, and perform attribute and structure reconstruction in augmented graphs. Innovatively, we coordinate the fusion of GMAEs with multiple masking mechanisms across multiplex graphs to reconstruct node attributes, network structures, and subgraphs in the original and augmented graphs. Furthermore, we learn to optimize node attributes and structural features through contrastive learning between original-view and augmented-view graphs to improve the model’s ability to detect anomalies. Meanwhile, we also propose a new anomaly score threshold selection strategy, which allows the model to be independent of the ground truth in real unsupervised scenarios. Extensive experiments on six real-world datasets with injected and real anomalies have demonstrated the effectiveness and efficiency of UMGAD compared with state-of-the-art UGAD methods.
@@
  • We propose a novel unsupervised multiplex graph anomaly detection method named UMGAD that emphasizes the importance of constructing and exploiting different interactions between nodes and addresses the UGAD problem in multiplex heterogeneous graphs.
@@
  • We conduct extensive experiments on datasets with both injected and real anomalies to demonstrate the superiority and efficiency of our proposed method. Experiment results show that UMGAD achieves 12.25%, and 11.29% average improvement in terms of AUC and Macro-F1 across six datasets compared to SOTA baselines.
@@ II Related Work
UGAD has gained significant attention recently, with many researchers focusing on using self-supervised signals to enhance anomaly detection in unsupervised settings. Early methods like Radar [27] and ResGCN [37] capture anomalies by analyzing the residuals of attribute information and its consistency with network data. With the rise of GNNs, UGAD models based on GNNs have emerged. We categorize GNN-based models into four types: MPI methods, CL-based methods, GAE-based methods, and MV methods.
@@
MPI methods. Most GNN-based methods adopt message passing [38] to learn node representations by aggregating neighbor information. However, inconsistencies between normal and anomalous features can hinder detection. To address this, Luo et al.[28] enhance message passing via community segmentation and structure learning. Bei et al.[39] employ reinforcement learning to expand neighbor pools, assess reliability, and amplify messages from trustworthy nodes. Qiao et al. [29] propose TAM, which maximizes local affinity on a truncated graph to avoid full-graph noise.
@@
CL-based methods. Contrastive learning, a popular approach for extracting self-supervised signals [40, 41, 42], has been applied to UGAD tasks [43] to capture feature inconsistencies between normal and anomalous nodes. CoLA [44] samples instance pairs using local network information to embed high-dimensional attributes and structures. Zhang et al.[45] introduce Sub-CR, which combines multi-view contrastive learning with attribute reconstruction. GRADATE[31] extends contrastive learning to the subgraph level via node-subgraph and subgraph-subgraph comparisons. Huang et al. [32] integrate variance-based learning with attribute reconstruction for balanced anomaly detection.
@@
Figure 1: The overview of the proposed UMGAD. The multiplex heterogeneous graph on the left takes Amazon dataset as an example, where U-P-U links users reviewing at least one product, U-S-U links users having at least one identical star rating within a week, and U-V-U links linguistically similar users.
@@
GAE-based methods. GAE-based approaches [46] detect anomalies by reconstructing node attributes and structural features. Ding et al.[47] use GCNs to capture graph topology and node attributes for embedding learning. Roy et al.[48] propose GAD-NR, which reconstructs a node’s neighborhood—including structure, self-properties, and neighbor features—based on its embedding. ADA-GAD [33] introduces a two-stage denoising framework: the GAE is first trained on denoised graphs, then the encoder is retrained on original graphs with anomaly regularization to avoid overfitting.
@@
MV methods. Recent work has explored anomaly detection on multiplex graphs, categorized as multi-view GAD methods. AnomMAN [49] detects anomalies in multi-view attribute networks by modeling cross-view attributes and interactions, using attention to fuse their importance. DualGAD [13] employs a dual-bootstrapped self-supervised framework to address feature-structure inconsistency and imbalanced anomaly signals. It integrates subgraph reconstruction with cluster-guided contrastive learning, achieving state-of-the-art performance in node-level anomaly detection.
Our UMGAD combines multi-view graph reconstruction with cross-view contrastive learning. Unlike single-view methods (e.g., GAE-based), it uses attribute-level and subgraph-level augmentations to capture richer relations. To avoid over-smoothing seen in MPI methods, UMGAD employs masking and GMAE fusion for selective feature focus. It also introduces a novel, label-free thresholding strategy for unsupervised anomaly detection.
@@ IV-A1 Attribute Reconstruction
where the original node attributes in $\mathcal{X}_{ma}$ are replaced by the [MASK] tokens, which are learnable vectors. Masked nodes are selected by utilizing uniform random sampling without replacement, which helps prevent potential bias. We repeat the attribute masking strategy on each relational subgraph for $K$ times to finally generate a set of masked subgraphs $\mathcal{G}_{ma}^{r}=\{\mathcal{G}_{ma}^{r,1},\mathcal{G}_{ma}^{r,2},\ldots,\mathcal{G}_{ma}^{r,K}\}$ for $\mathcal{G}^{r}$. These masked subgraphs are then fed into the GNN-based encoder and decoder, and the output of each encoder is:
@@
where $\mathcal{X}_{ma}^{r,k}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the output of the attribute decoder in the $k$-th masking repeat of the $r$-th relational subgraph. $\mathcal{V}_{re}=\mathcal{V}\setminus\mathcal{V}_{ma}$ denotes the set of remaining nodes that are unaffected by masking, $\mathbf{W}_{enc1}^{r,k}\in\mathbb{R}^{f\times d_{h}}$ and $\mathbf{W}_{dec1}^{r,k}\in\mathbb{R}^{d_{h}\times f}$ are the trainable weight matrices, $d_{h}$ represents the dimension of latent vectors during graph convolution. Inside each encoder and decoder, before the vertical lines |||| are the input variables, and after the vertical lines are the important learnable weights. In subsequent formulas, each vertical line means the same.
After processing each decoupled relational subgraph, UMGAD learns the importance (weights) of different relations collaboratively, instead of dealing with each relational subgraph independently. It fuses the effects of these relations on anomaly detection through a set of learnable weight parameters $a^{r}$, learnable weights as follows:
@@ IV-A2 Structure Reconstruction
where $g(v,u)={x_{ms}^{r,k}(v)}^{\mathsf{T}}x_{ms}^{r,k}(u)$ calculates the estimated probability of link between node $v$ and $u$, and then we can obtain the reconstructed adjacent matrix $\tilde{\mathbf{A}}_{O}^{r}$. We introduce negative sampling to train a more generalized model by accumulating all edge prediction probabilities of the unmasked subgraphs in the denominator, making the model robust to external noise. Notably, the standard normalization in Eq. (4) is used to calculate the similarity between the reconstructed attribute features and the original attribute features. The softmax normalization in Eq. (7) is used to evaluate the probability of a link being reconstructed between two nodes.
Similar to Eq. (3), collaboratively considering the importance of different relational subgraphs for structure reconstruction, we aggregate all subgraph structure reconstruction losses together using a set of learnable weight parameters $b^{r}$ as follows:
@@ IV-B2 Subgraph-level Augmented Graph Reconstruction
Subgraph-level augmentation can be considered as a specific combination of node attribute level and structure level augmentation. The reconstructed node attribute matrix $\tilde{\mathcal{X}}_{sa}^{k}$ and structural adjacency matrix $\tilde{\mathbf{A}}_{ss}^{r}$ are shown below, respectively:
@@
Then the subgraph-level augmented attribute and structure reconstruction loss values $\mathcal{L}_{sa}$ and $\mathcal{L}_{ss}$ are defined as follows:
@@
It is worthwhile to emphasize that normal nodes tend to have more similar features than their neighbors, while anomalous nodes exhibit greater feature differences. Through masking and GNN-based reconstruction, normal nodes are more likely to be accurately reconstructed, whereas anomalous nodes are less likely to be. Therefore, we improve the model’s ability to identify anomalies by employing different masking mechanisms with a graph autoencoder.
@@ IV-C Dual-view Contrastive Learning
Dual-view contrastive learning is defined between the original view and the augmented view to learn more representative and intrinsic node embeddings, which helps capture further anomaly information. Specifically, we optimize the node attribute under joint loss by comparing the original-view graph with the attribute-level augmented-view graph and the subgraph-level augmented-view graph, respectively. The node attribute of node $i$ in the original view forms a positive pair with the augmented view and a negative pair with the attribute of node $j$ in the original and augmented views. We use the following loss function to optimize contrastive learning:
@@ IV-D Optimization Objective
Putting it all together, we have the The overall loss function in the training stage is as follows:
@@
where {*} denotes the specific original $(O)$, attribute-level augmented $(A\_Aug)$, or subgraph-level augmented $(S\_Aug)$ graph, $||\cdot|{|_{1}}$ denotes the Euclidean norm of a vector, $||\cdot|{|_{2}}$ denotes the L1-norm of the vector, $\tilde{\zeta}_{*}^{r}(i)$ and $\zeta^{r}(i)$ represent the $i$-th row of the specific reconstructed structure matrix $\tilde{\mathbf{A}}_{*}^{r}$ and original structure matrix $\mathbf{A}^{r}$, respectively. Similarly, $\tilde{x}_{*}(i)$ and $x(i)$ are the reconstructed attribute vector and original attribute vector of the node $i$, respectively. Therefore, the anomaly score $S(i)$ of node $i$ is the mean value of all $S(i)_{*}$.